# Bridging Model Reference Adaptive Control and Data Informativity

### v2

@@ 6 Simulation Results
[…]
@@ 6.1 Numerical system
(a) State tracking errors in $\mathcal{S}_{1}$
(b) Norm of matching error in $\mathcal{S}_{1}$
(c) Input signals in $\mathcal{S}_{1}$
(d) State tracking errors in $\mathcal{S}_{2}$
(e) Norm of matching error in $\mathcal{S}_{2}$
(f) Input signals in $\mathcal{S}_{2}$
@@ 6.2 Highly maneuverable aircraft example
We consider the unstable longitudinal dynamics of a highly maneuverable aircraft [hartmann1979control, yuan2018robust] with three states (angle of attack, pitch rate, and pitch angle) and four inputs (elevator, elevon, canard, and symmetric aileron). For better numerical conditioning, we scale the angle of attack by 100. With sampling time of 0.01s, we get a system model in the form (1), with
@@
The choice $B_{\rm m}=B_{\rm s}$ is based on the setting in [hartmann1979control, yuan2018robust]. Nevertheless, we stress on the fact that $B_{\rm s}$ is unknown: hence, even though we have $B_{\rm m}=B_{\rm s}$ in the simulation, we do not know that $B_{\rm m}=B_{\rm s}$ for control design.
@@
(a) State tracking errors in $\mathcal{S}_{5}$
(b) Norm of matching error in $\mathcal{S}_{5}$
(c) Input signals in $\mathcal{S}_{5}$
(d) State tracking errors in $\mathcal{S}_{6}$
(e) Norm of matching error in $\mathcal{S}_{6}$
(f) Input signals in $\mathcal{S}_{6}$
@@
(a) State tracking errors in $\mathcal{S}_{7}$
(b) Norm of matching error in $\mathcal{S}_{7}$
(c) Input signals in $\mathcal{S}_{7}$
(d) State tracking errors in $\mathcal{S}_{8}$
(e) Norm of matching error in $\mathcal{S}_{8}$
(f) Input signals in $\mathcal{S}_{8}$
@@ 7 Conclusions
In this paper, the MRAC problem has been investigated from a novel perspective of data informativity. Different from previous results on data informativity that used data generated offline, in this paper we have taken an online perspective by verifying data informativity at every time step. A new MRAC method has been developed that ensures asymptotic convergence of the adaptive gains to a solution of the matching equations, provided such a solution exists. As a preliminary problem, we have analyzed the relationship between data informativity for (unique) system identification and data informativity for model reference control, demonstrating that the former is only sufficient, but not necessary for informativity for model reference control. It becomes necessary only in specific scenarios. Motivated by these findings, we have devised an input function that switches between an adaptive controller and a term that increases the rank of the data matrix. Under the assumption that the matching equations have a solution, this input generates informative data for model reference control after a finite number of time steps. Moreover, in this situation, the adaptive gains of the controller are shown to converge to a solution of the matching equations. As compared to state-of-the-art MRAC approaches, the proposed method does not need knowledge of the input matrix and does not require uniqueness of the solution of the matching equations. A possibility for future research is to explore partial-information scenarios where only input and output data are available.

### v3

@@ 6 Simulation Results
with $\varepsilon=10^{-10}$. During the implementation, the informative time $T^{*}$ is initialized to $T^{*}=\infty$. At every time step $t$, we check whether the data $(U_{-}(t),X(t))$ satisfy (11), and update $T^{*}$ accordingly. In all simulations, we verify that $T^{*}\leq n+m$, consistent with Lemma 5.9, and that the control gains $(\hat{K}(t),\hat{L}(t))$ in (29) converge to a solution of the matching equations (6), as guaranteed by Theorem 5.14.
@@ 6.1 Numerical system
(a) State tracking errors in $\mathcal{S}_{1}$
(b) Norm of matching error in $\mathcal{S}_{1}$
(c) Input signals in $\mathcal{S}_{1}$
(d) State tracking errors in $\mathcal{S}_{2}$
(e) Norm of matching error in $\mathcal{S}_{2}$
(f) Input signals in $\mathcal{S}_{2}$
@@ 6.2 Highly maneuverable aircraft example
We consider the unstable longitudinal dynamics of a highly maneuverable aircraft [17, 44] with three states (angle of attack, pitch rate, and pitch angle) and four inputs (elevator, elevon, canard, and symmetric aileron). For better numerical conditioning, we scale the angle of attack by 100. With sampling time of 0.01s, we get a system model in the form (1), with
@@
The choice $B_{\rm m}=B_{\rm s}$ is based on the setting in [17, 44]. Nevertheless, we stress on the fact that $B_{\rm s}$ is unknown: hence, even though we have $B_{\rm m}=B_{\rm s}$ in the simulation, we do not know that $B_{\rm m}=B_{\rm s}$ for control design.
(a) State tracking errors in $\mathcal{S}_{5}$
(b) Norm of matching error in $\mathcal{S}_{5}$
(c) Input signals in $\mathcal{S}_{5}$
(d) State tracking errors in $\mathcal{S}_{6}$
(e) Norm of matching error in $\mathcal{S}_{6}$
(f) Input signals in $\mathcal{S}_{6}$
@@
(a) State tracking errors in $\mathcal{S}_{7}$
(b) Norm of matching error in $\mathcal{S}_{7}$
(c) Input signals in $\mathcal{S}_{7}$
(d) State tracking errors in $\mathcal{S}_{8}$
(e) Norm of matching error in $\mathcal{S}_{8}$
(f) Input signals in $\mathcal{S}_{8}$
@@ 7 Conclusions
In this paper, the MRAC problem has been investigated from a novel perspective of data informativity. Different from previous results on data informativity that used data generated offline, in this paper we have taken an online perspective by verifying data informativity at every time step. A new MRAC method has been developed that ensures asymptotic convergence of the adaptive gains to a solution of the matching equations, provided such a solution exists. As a preliminary problem, we have analyzed the relationship between data informativity for (unique) system identification and data informativity for model reference control, demonstrating that the former is only sufficient, but not necessary for informativity for model reference control. It becomes necessary only in specific scenarios. Motivated by these findings, we have studied necessary and sufficient conditions for the existence of an adaptive law that guarantees convergence of the adaptive gains to a solution of the matching equations. Based on this existence analysis, we have devised an input function that switches between an adaptive controller and a term that increases the rank of the data matrix. Under the assumption that the matching equations have a solution, this input generates informative data for model reference control after a finite number of time steps. Moreover, in this situation, the adaptive gains of the controller are shown to converge to a solution of the matching equations. As compared to state-of-the-art MRAC approaches, the proposed method does not need knowledge of the input matrix and does not require uniqueness of the solution of the matching equations. A possibility for future research is to explore partial-information scenarios where only input and output data are available.